OpenStudy (anonymous):

Algebra 2 Trouble :/

OpenStudy (anonymous):

no

OpenStudy (anonymous):

In answer A, the equation for the line -x-2 is incorrect. Remember that y=-x-2 has a slope of negative 1. You can see in answer A that the line has a positive slope, so it's not the right line

OpenStudy (anonymous):

\(y=-x-2\) is a line with a negative slope (namely \(-1\)) there is only one choice with a line of negative slope

OpenStudy (anonymous):

Ok, well if you have an equation of a line like y=mx+b, where m and b are numbers, the number m is the slope. And a positive slope line points up and to the right, and a negative slope line points up and to the left

OpenStudy (anonymous):



OpenStudy (anonymous):

Yep, and since it's the only graph where the line has a negative slope you know it has to be the right answer

OpenStudy (anonymous):

C, no doubt about it! Yes! Wow, surprisingly I actually learned something I know I will remember!

OpenStudy (anonymous):

Good to hear! Feel free to tag me in any other questions you have

OpenStudy (anonymous):

@vinnv226 Is there any way I can give you like a double boost or some kind of extra recognition on here? Because you truly deserve it, man.

OpenStudy (anonymous):

Thanks, I'm just glad I was able to help you. You could leave me a testimonial on my profile if you'd like to

OpenStudy (anonymous):

That's a done deal!